OpenStudy (anonymous):

A bucket that contains several inches of water is left outside on a rainy day. The height of the water in the bucket that day is given by the function equation H = 4 + 0.5t, where H is the height of the water and t is the time in hours the bucket sits out in the rain.What is the meaning of the vertical intercept in this equation? A.The bucket originally contained 0.5 in. of water. B.The bucket originally contained 4 in. of water. C.The bucket fills at a rate of 0.5 in./h. D.The bucket fills at a rate of 4 in./h.

OpenStudy (anonymous):

@AndrewTheCookie @huntin14 @Gokuporter @lexie<3 @lexismith @ObEy-grd-newsome @One098 @Pawtpie @pooja195

OpenStudy (anonymous):

The bucket originally contained 4 in. of water.

OpenStudy (blake57roger):

b

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

np

OpenStudy (blake57roger):

an equation is y = mx + b and the Y intercept is B so if we reform the equation: H = 4 + 0.5t into H = 0.5t + 4 then 4 is the Y intercept

OpenStudy (blake57roger):

medal

OpenStudy (anonymous):

nope u copied from the internet no cheaters get medals